# E2ENet: Dynamic Sparse Feature Fusion for Accurate and Efficient 3D Medical Image Segmentation

### v1

@@ 4.5.1 Effect of DSFF Mechanism
4.5.2 Effect of Restricted Depth-Shift in 3D Convolution
We evaluated the effectiveness of our restricted shift strategy by comparing the performance of E2ENet and its variants. Without restricted depth-shift, the mDice decreased from 74.5% to 73.9% for the BraTS challenge as shown in 3rd row of Table 4, and decreased from 90.1% to 88.2% for AMOS-CT (as shown in the 3rd row of Table 5). In Table 5, when comparing E2ENet with kernel sizes of 3x3x3 (last 3rd and 4th row) to E2ENet with kernel sizes of 1x3x3 with a depth shift (2nd and 6th row), their segmentation accuracy remains the same, whether with DSFF or without DSFF. This demonstrates that the use of a 1x3x3 kernel with restricted depth shift is functionally equivalent to a 3x3x3 kernel in terms of segmentation accuracy, all the while offering savings in computational and memory resources.
Rows 7 and 8 of Table 4 demonstrate that when restricted shift is applied to the height or width dimensions, the model’s performance decreases compared to E2ENet (4th row) with restricted shift on the depth dimension, called as restricted depth-shift. We also observed that the performance of 3D convolution with kernel sizes of $1\times 3\times 3$ (row 3), $3\times 1\times 3$ (row 5), and $3\times 3\times 1$ (row 6) decreased significantly without restricted shift compared to their counterparts with restricted shift on the corresponding dimensions (rows 4, 7, and 8, respectively). Furthermore, Table 5 shows that in the AMOS-CT challenge, models without restricted depth-shift (1st and 3rd row) exhibited a noticeable decline in performance compared to their counterparts with restricted depth-shift. These phenomena demonstrate the effectiveness of restricted shift, especially restricted depth-shift, for medical image segmentation.
Furthermore, we evaluated the impact of different shift sizes on the model’s performance. As shown in Table 6, we compared the performance of E2ENet with shift sizes of (-1, 0, 1) to (-2, 0, 2) and (-3, 0, 3), and observed a decrease in performance from $90.1\%$ to $89.8\%$ and $89.7\%$, respectively, as the shift size increased. This demonstrates that excessive shifting can lead to the overrepresentation of depth information in sparse feature maps, which can negatively impact the continuity of 3D image information. Therefore, we use a shift size of (-1, 0, 1) in our restricted depth-shift strategy as the default setting in our experiments.
Finally, to demonstrate the advantages of individual modules, we plot a critical distance diagram using the Nemenyi post-hoc test with a p-value of 0.05 to establish the statistical significance of our modules. In Figure 6, the top line represents the axis along which the methods’ average ranks, and a lower value indicates better performance. Methods joined by thick horizontal black lines are considered not statistically different. From the diagram, we can clearly observe that E2ENet with depth shift significantly outperforms E2ENet without depth shift. Additionally, the incorporation of dynamic sparse feature fusion into E2ENet results in a substantial reduction in both the number of FLOPs (from 23.90M to 11.23M) and parameters (from 3069.55G to 969.32G) while maintaining comparable performance, without any significant performance degradation.
4.6 The Impact of Weighting Factors $\alpha$ on PT Score
@@
4.7 Convergence Analysis
@@
4.8 Generalizability Analysis
Table 7: Evaluating the generalizability of E2ENet on AMOS-MRI dataset, CT $\rightarrow$ MRI: using CT dataset for pretraining and MRI dataset for fine-tuning; MRI: using only MRI dataset for training; CT+MRI: using both CT and MRI dataset for training. The results, except for E2ENet and nnUNet (CT $\rightarrow$ MRI), are collected from the AMOS website.
method mDice mNSD
CT $\rightarrow$ MRI MRI CT + MRI CT $\rightarrow$ MRI MRI CT + MRI
E2ENet 87.8 / / 83.9 / /
nnUNet 87.4 87.1 87.7 83.3 83.1 82.7
VNet / 83.9 75.4 / 65.8 65.6
nnFormer / 80.6 75.48 / 74.0 66.63
CoTr / 77.5 73.46 / 78.0 65.35
Swin UNTER / 75.7 77.52 / 65.3 69.10
UNTER / 75.3 / / 70.1 /
Table 8: Quantitative comparisons of segmentation performance on AMOS-CT test dataset. The mDice and mNSD of these baselines, except for nnUNet, are collected from the AMOS challenge (Ji et al., 2022). ${}^{{\dagger}}$ indicates the results without postprocessing that are reproduced by us. ${}^{{\ddagger}}$ denotes the results with postprocessing that are reproduced by us. $*{}^{*}$ indicates the results with postprocessing.
Method mDice Params FLOPs 1 PT score-Test mNSD
CoTr $80.9$ 41.87 1510.53 1.11 66.3
nnFormer $85.6$ 150.14 1343.65 1.11 72.5
UNETR $79.4$ 93.02 391.03 1.41 60.8
Swin-UNETR $86.3$ 62.83 1562.99 1.13 73.8
VNet $82.9$ 45.65 1737.57 1.11 67.6
${}^{{\dagger}}$ 90.6 30.76 1067.89 1.31 82.0
${}^{{\ddagger}}$ 91.0 30.76 1067.89 1.31 82.6
$*{}^{*}$ (s=0.7) 90.7 11.23 969.32 1.54 82.2
$*{}^{*}$ (s=0.8) 90.7 9.44 778.74 1.65 82.1
$*{}^{*}$ (s=0.9) 90.4 7.64 492.29 1.89 81.4
E2ENet (s=0.7) 90.6 11.23 969.32 1.54 82.0
E2ENet (s=0.8) 90.4 9.44 778.74 1.65 81.8
E2ENet (s=0.9) 90.1 7.64 492.29 1.89 80.7
  • 1
    The inference FLOPs is calculated based on the patch sizes of $1\times 128\times 128\times 128$ without considering postprocessing cost.
4.9 Model Capacity Analysis
To ensure a fair comparison, we adjusted the configuration of nnUNet by reducing its width from 32 to 27 and its depth from 5 to 4, resulting in a modified version referred to as nnUNet (-). Additionally, we enhanced E2ENet by increasing its width from 48 to 58, creating a modified version referred to as E2ENet(+). We evaluated the performance of these models, and the results can be found in Table 9. It is worth noting that scaling down nnUNet resulted in decreased performance in terms of mDice and mNSD, while scaling up E2ENet led to an increase in mDice and comparable performance in terms of mNSD. This indicates that the comparable performance of the memory and computation-efficient E2ENet is not attributed to the dataset’s requirement for a small number of parameters and computations.
Table 9: A comparison under a similar FLOPs budget on the AMOS-CT dataset in both cases: when the nnUNet is scaled down, or E2ENet is scaled up.
Method mDice Params FLOPs 1 mNSD
E2ENet (S=0.8) 90.0 9.43 778.74 82.3
nnUNet 90.0 31.20 1067.89 82.1
nnUNet (-) 89.7 12.96 755.79 81.9
E2ENet (+) 90.4 10.37 1119.17 82.2
  • 1
    The inference FLOPs is calculated based on the patch sizes of $1\times 128\times 128\times 128$ without considering postprocessing cost.
4.9.1 BraTS Challenge in MSD
4.10 Qualitative Results
In this section, we compare the proposed E2ENet and nnUNet qualitatively on three challenges. To make the comparison easier, we highlight detailed segmentations with red dashed boxes.
4.10.1 AMOS-CT Challenge
Figure 10 (a) presents a qualitative comparison of our proposed E2ENet with nnUNet on the AMOS-CT challenge. As a widely used baseline model, nnUNet has been evaluated on multiple medical image segmentation challenges. Our results demonstrate that, on certain samples, E2ENet can improve segmentation quality and overcome some of the challenges faced in the AMOS-CT challenge. For example, as shown in the first column, E2ENet accurately distinguishes between the stomach and esophagus, which can be a difficult task due to their close proximity. In the second column, E2ENet better differentiates the duodenum from the background, while in the third column, E2ENet accurately identifies the precise boundaries of the liver, a structure that is prone to over-segmentation. These examples demonstrate the potential of our proposed method to improve the accuracy of medical image segmentation to some extent, especially in challenging cases such as distinguishing between closely located organs or accurately segmenting complex shapes.
4.10.2 BTCV Challenge
In Figure 10 (b), we present a qualitative comparison of our proposed E2ENet method with nnUNet as a baseline model on the BTCV challenge. Our results demonstrate the effectiveness of our proposed method in addressing some of the challenges of medical image segmentation. For example, as shown in the first and third columns, our E2ENet method accurately distinguishes the stomach from the background without over- or under-segmentation, which can be difficult due to the low contrast in the image. In the second column, E2ENet performs well in differentiating the stomach from the spleen. These examples suggest that our DSFF module can effectively encode feature information for improved performance in medical image segmentation.
Figure 11 presents a qualitative comparison of our proposed E2ENet method with the nnUNet on the BraTS challenge with highly variable shapes of the segmentation targets. Based on the results of the baseline model, nnUNet, we observed that accurately distinguishing the edema (ED) from the background is difficult, as the edema tends to have less smooth boundaries. Our results suggest that E2ENet may have some potential to improve the distinguishability of the edema boundaries, as evidenced by the relatively better segmentation results in the first, second, and fourth columns. Moreover, E2ENet accurately differentiates the enhanced tumor (ET) from the edema, as shown in the third column, which is a challenging task due to the similarity in appearance between these two regions, and the dispersive distribution of ET. These findings suggest that E2ENet is a promising method for accurately segmenting brain tumors in challenging scenarios.
4.11 Feature Fusion Visualization
Figure 12: The proportions of feature map connections before and after training with the DSFF mechanism at a sparsity level of 0.8 on AMOS-CT, BTCV, and BraTS challenges. The proportions before training and the proportions after training are shown in (a), (b), and (c) for the three challenges respectively.
Figure 13: The proportions of feature map connections during training with the DSFF mechanism at a sparsity level of 0.8 on AMOS-CT.
In this section, we explore how the DSFF mechanism fuses features from three directions (upward, forward, and downward) during training. To shed light on this mechanism, we specifically analyze the proportions of feature map connections from different directions to a specific fused feature node, considering a feature sparsity level of 0.8. Figure 12 provides a visualization of these proportions obtained after training on the challenging AMOS, BTCV, and BraTS datasets, and Figure 13 provides a visualization of these proportions obtained during training on the challenging AMOS. We calculated these proportions by dividing the number of non-zero connections in a given direction by the total number of non-zero connections within the fused feature node.
After training, we observed in Figure 12 (b)-(d) that the DSFF module learned to assign greater importance to features from the ”forward” directions for most fused feature nodes. For example, in the first feature level, the flow and processing of features can be observed as if they were passing through a fully convolutional neural network (FCN), which preserves the spatial dimensions of the input image. At the second feature level, the original image is downsampled by a factor of $1/2$ and flows through another FCN. Therefore, from this perspective, E2ENet can take the advantage of FCN effectively incorporate and preserve multi-scale information, while reducing computation cost.
Simultaneously, the complementary feature flows from the ”upward” and ”downward” directions provide richer information. At the low feature levels and the earlier fusion step, the “upward” information is more dominant than the “downward” information. This prioritization of upward feature flow information is similar to the design of the decoder in UNet and UNet++, while alleviating the semantic gap. As the levels and fused steps increase, the proportion of feature map connections in the “downward” direction increases, which allows the network to effectively preserve low-level information and integrate it with high-level information, leading to improved performance in capturing fine details. Interestingly, this trend becomes increasingly apparent during training, as illustrated in Figure 13.
4.12 Organ Volume Statistics and Class-wise Results Visualization
@@
Figure 14: (a) The organ volume statistics of AMOS-CT training dataset. (b) Class-wise Dice of nnUNet without postprocessing (visualization of Table 1). (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ without postprocessing and nnUNet without postprocessing on AMOS-CT validation dataset . The positive value means that E2ENet outperforms nnUNet, vice versa.)
@@
Figure 15: (a) The organ volume statistics of BTCV training dataset. (b) Class-wise Dice of nnUNet (visualization of Table 2). Note that AG denotes the average of the right and left adrenal glands (RAG and LAG). (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ and nnUNet on BTCV test dataset. The positive value means that E2ENet outperforms nnUNet, vice versa.
@@
Figure 16: (a) The organ volume statistics of BraTS training dataset. (b) Class-wise Dice of nnUNet (visualization of Table 3) (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ and nnUNet on 5-fold cross-validation of the training dataset. The positive value means that E2ENet outperforms nnUNet, vice versa.
@@
In this section, we analyzed the relationship between organ volume and segmentation accuracy on the AMOS-CT, BTCV, and BraTS challenges. The results, depicted in Figures 14, 15 and 16, showed that small organs with relatively low segmentation accuracy. For the AMOS-CT challenge, RAG (right adrenal gland), LAG (left adrenal gland), Gall (gallbladder), and Eso (esophagus) are more challenging to accurately segment. This may be due to the fact that smaller organ volumes provide less visual information for the segmentation algorithm to work with. However, our proposed method, E2ENet, also demonstrated comparable (or better) performance on these small organs, particularly for the organ “LAG”, in which the Dice improved from $81.7\%$ to $82.4\%$. On the BTCV challenge, the Dice of “Gall”, which is considered to be the most challenging organ, improves from $75.3\%$ to $78.1\%$ when using E2ENet compared to nnUNet. For the BraTs challenge, E2ENet demonstrates the most significant improvement in the Dice score of the ”ET” region, which is considered the most challenging class, with an increase of $0.7\%$.
@@
5 Conclusion
In this work, our aim is to address the challenge of designing a 3D medical image segmentation method that is both accurate and efficient. By proposing a dynamic sparse feature fusion mechanism and incorporating restricted depth-shift in 3D convolution, our proposed E2ENet model is able to improve performance on 3D medical image segmentation tasks while significantly reducing the computation and memory overhead. The dynamic sparse feature fusion mechanism has shown that it adaptively learns the importance of each feature map and zeros out the less important ones. This leads to a more efficient feature representation without sacrificing performance. Additionally, the experiments demonstrate that the restricted depth-shift in 3D convolution enables the model to capture spatial information more effectively and efficiently.
Extensive experiments on three benchmarks show that E2ENet consistently achieves a superior trade-off between accuracy and efficiency compared to prior state-of-the-art baselines. While E2ENet provides a promising solution for balancing accuracy and computational cost, future work could explore the potential of learnable shift offset, which may lead to even better performance. Additionally, the effectiveness of E2ENet strongly depends on the choice of the initial architecture. It is interesting to apply the DSFF modules and shift convolution to different architectures to explore their potential in the future. Furthermore, future advancements in hardware support for sparse neural networks could fully unlock the potential of sparse training methods across further resource-constrained applications.
6 Acknowledgement
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-4456 and EINF-5565.

### v2

@@ A.8.3 Statistical Significance of Designed Modules
To demonstrate the advantages of individual modules, we plot a critical distance diagram using the Nemenyi post-hoc test with a p-value of 0.05 to establish the statistical significance of our modules. In Figure 7, the top line represents the axis along which the methods’ average ranks, and a lower value indicates better performance. Methods joined by thick horizontal black lines are considered not statistically different. From the diagram, we can clearly observe that E2ENet with depth shift significantly outperforms E2ENet without depth shift. Additionally, the incorporation of dynamic sparse feature fusion into E2ENet results in a substantial reduction in both the number of FLOPs (from 23.90M to 11.23M) and parameters (from 3069.55G to 969.32G) while maintaining comparable performance, without any significant performance degradation.
A.9 The Impact of Weighting Factors $\alpha$ on PT Score
@@
A.9.1 Qualitative Results
BTCV Challenge In Figure 9 (b), we present a qualitative comparison of our proposed E2ENet method with nnUNet as a baseline model on the BTCV challenge. Our results demonstrate the effectiveness of our proposed method in addressing some of the challenges of medical image segmentation. For example, as shown in the first and third columns, our E2ENet method accurately distinguishes the stomach from the background without over- or under-segmentation, which can be difficult due to the low contrast in the image. In the second column, E2ENet performs well in differentiating the stomach from the spleen. These examples suggest that our DSFF module can effectively encode feature information for improved performance in medical image segmentation.
BraTS Challenge in MSD Figure 10 presents a qualitative comparison of our proposed E2ENet method with the nnUNet on the BraTS challenge with highly variable shapes of the segmentation targets. Based on the results of the baseline model, nnUNet, we observed that accurately distinguishing the edema (ED) from the background is difficult, as the edema tends to have less smooth boundaries. Our results suggest that E2ENet may have some potential to improve the distinguishability of the edema boundaries, as evidenced by the relatively better segmentation results in the first, second, and fourth columns. Moreover, E2ENet accurately differentiates the enhanced tumor (ET) from the edema, as shown in the third column, which is a challenging task due to the similarity in appearance between these two regions, and the dispersive distribution of ET. These findings suggest that E2ENet is a promising method for accurately segmenting brain tumors in challenging scenarios.
A.10 Convergence Analysis
@@
A.11 Organ Volume Statistics and Class-wise Results Visualization
@@
Figure 13: (a) The organ volume statistics of AMOS-CT training dataset. (b) Class-wise Dice of nnUNet without postprocessing (visualization of Table 1). (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ without postprocessing and nnUNet without postprocessing on AMOS-CT validation dataset. The positive value means that E2ENet outperforms nnUNet, and vice versa.)
@@
Figure 14: (a) The organ volume statistics of BTCV training dataset. (b) Class-wise Dice of nnUNet (visualization of Table 11). Note that AG denotes the average of the right and left adrenal glands (RAG and LAG). (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ and nnUNet on BTCV test dataset. The positive value means that E2ENet outperforms nnUNet, and vice versa.
@@
Figure 15: (a) The organ volume statistics of BraTS training dataset. (b) Class-wise Dice of nnUNet (visualization of Table 2) (c) Class-wise Dice differences between E2ENet with feature sparsity $0.7$ and nnUNet on 5-fold cross-validation of the training dataset. The positive value means that E2ENet outperforms nnUNet, and vice versa.
@@
In this section, we analyzed the relationship between organ volume and segmentation accuracy on the AMOS-CT, BTCV, and BraTS challenges. The results, depicted in Figures 13, 14 and 15, showed that small organs with relatively low segmentation accuracy. For the AMOS-CT challenge, RAG (right adrenal gland), LAG (left adrenal gland), Gall (gallbladder), and Eso (esophagus) are more challenging to accurately segment. This may be due to the fact that smaller organ volumes provide less visual information for the segmentation algorithm to work with. However, our proposed method, E2ENet, also demonstrated comparable (or better) performance on these small organs, particularly for the organ “LAG”, in which the Dice improved from $81.7\%$ to $82.4\%$. On the BTCV challenge, the Dice of “Gall”, which is considered to be the most challenging organ, improves from $75.3\%$ to $78.1\%$ when using E2ENet compared to nnUNet. For the BraTs challenge, E2ENet demonstrates the most significant improvement in the Dice score of the "ET" region, which is considered the most challenging class, with an increase of $0.7\%$.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have claimed that our proposed Efficient to Efficient Network (E2ENet) in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The efficiency in this paper is theoretical, we still need future work for hard work support, which we claim in conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [No]
  14. Justification: We don’t include theoretical results for proof
  15. 4.
    Experimental Result Reproducibility
  16. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  17. Answer: [Yes]
  18. Justification: We provide implementation details in A.5, and the code and dataset links in the supplementary materials.
  19. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  20. 5.
    Open access to data and code
  21. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  22. Answer: [Yes]
  23. Justification: The dataset we use is accessible online, and the code is provided in the supplementary materials.
  24. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: We provide implementation details in A.5.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [Yes]
  33. Justification: In Section A.8.3, we plot a critical distance diagram using the Nemeny post-hoc test with a p-value of 0.05 to establish the statistical significance of our modules.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: We provide the type of hardware used for the implementation in Appendix A.5.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: We have reviewed the NeurIPS Code of Ethics.
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: We include border impact in Section A.1.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: The paper poses no such risks.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: The Licenses names for existing assets are CC-BY 4.0.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: Our paper does not release new assets.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.